OpenStudy (anonymous):

simplify (4x^5y^3z)(5xy^6z^6)

OpenStudy (anonymous):

20x^6y^9z^7

OpenStudy (anonymous):

do you know... 22a^3b^5 all divided by 11ab^3c^2

OpenStudy (anonymous):

2a2b2/c2 i think im not sure

OpenStudy (anonymous):

2a^2b^2/c^2 think so,

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

anymore?

OpenStudy (anonymous):

the llength of a rectangle is (3x+6) and the width is (2x+6) find the area?

OpenStudy (anonymous):

k i think it's 6x^2+30x+36,

OpenStudy (anonymous):

fan me back

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

k anymore?

OpenStudy (anonymous):

okay this is the last two.... what are the solutions to the equation (x-2)(x+5)=0 what are the solutions to the equations X^2+9x+14=0

OpenStudy (anonymous):

im solving the equation for x right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the 1st one is x=2,−5 the second one is x=−7,−2

OpenStudy (anonymous):

if thats all of them, im glad I could help xD